OpenStudy (crystal_b16):

Help! My Algebra 1 teacher handed out a hand written assignment. I need some help on how to solve this equation: 5x – 5 >= 10 or -3x + 1 >= 13 She says I have to show all of my work. I'm not that good on solving these equations. It would help me in the second part of the assignment also. I need to graph the inequality in the second part *if you notice answers below, don't mind them. Their from a different question. I'm unsure of how to delete previous questions and ask new ones other then editing them and asking a new one* Can someone please help me? This is due by midnight

alones (alones):

I'll help you with the first one but you post each question per post oki?

OpenStudy (steve816):

oki!

OpenStudy (mathsucks321):

A. x<= -10 First one^

alones (alones):

Multiply each sides by \(2\), than simpltiply \(5\) by \(2\) to \(10\) \[\ 10≤−x\] And multiply each side by negative one \[\ −10≥x\] Than switch those side, \(x≤−10\) or as \(x<= -10\)

alones (alones):

*Simplify >_>

OpenStudy (crystal_b16):

Here. This file has both the equation and the graph. If your able to help, please do.

OpenStudy (kaylieanime):

The first half has already been helped with, unless you still don't understand it. The second half of the same question was over looked: \[-3x+1\ge13\] Would you like help with that?

OpenStudy (crystal_b16):

Yes please

OpenStudy (kaylieanime):

You must isolate x first. I'll do the first step. \[-3x+1\ge13\] \[-3x \ge 12\] Could you try to do the rest?

OpenStudy (crystal_b16):

I could try. But I'm not really sure what else to do in order to try

OpenStudy (kaylieanime):

You want x as one unit. You want x as x (it looks the same as 1x), not -3x.

OpenStudy (crystal_b16):

Would I want to divide both sides by -3 (if I have to divide) or one side by -3 and the other by positive 3?

OpenStudy (kaylieanime):

Both sides by -3. You will also flip the sign because the sign flips when either multiplying or dividing

OpenStudy (crystal_b16):

Being that -3 cancels itself out and i get -4 when i divide -3 by 12, how would I set up the inequality?

OpenStudy (kaylieanime):

Applying the rules (sign changing, x isolated...) it would look like \[x \le-4\] Did you understand when the others helped you with the 5x-5>= 10?

OpenStudy (crystal_b16):

Not really no. This concept is a bit confusing for me

OpenStudy (kaylieanime):

Ok this one looks easier. Add five to isolate x \[5x-5\ge10\] changes into \[5x \ge15\]

OpenStudy (kaylieanime):

Make sense?

OpenStudy (crystal_b16):

Yep

OpenStudy (kaylieanime):

The last step in turning 5x into x is to...?

OpenStudy (crystal_b16):

Cancel out 5? I'm honestly not really sure

OpenStudy (kaylieanime):

Divide all sides by 5. The end result would be...?

OpenStudy (crystal_b16):

X ≥ 3?

OpenStudy (kaylieanime):

Yes!

OpenStudy (kaylieanime):

The second half is visually putting the two inequalities together

OpenStudy (crystal_b16):

How exactly would I do that?

OpenStudy (kaylieanime):

By taking each inequality and graphing it on a number line. Let's take a finished one. \[x \ge5\] and plotting it on the line. The line should look like this: 

OpenStudy (kaylieanime):

The picture looks bad, but each tick represents a number with each side extending to infinity

OpenStudy (kaylieanime):

\[x \ge5\] would start on exactly 5 and stretch infinity positively. The line underneath the > means equal, so the symbol entirely would 'exactly 'x' (whatever number that might be) and above.' Is 6 bigger than 5? Yes, so we say \[6\ge5\] Understand?

OpenStudy (crystal_b16):

Yep

OpenStudy (crystal_b16):

Then the other one would extend in the opposite direction correct?

OpenStudy (kaylieanime):

Yes. \[x \le-4\] would strech the other way. You'll be seeing 'and' and 'or' a lot of times. Here are the rules. And: the x is in between the two inequalities Or: the x is either in one inequality or the other but not both.

OpenStudy (kaylieanime):

In this case, it's an 'or' so the unknown x is in one inequality. This will show a gap REMEMBER: In inequalities there is no definite number for x hence the IN in inequality. How would the number line look like with the inequalities shown?

OpenStudy (crystal_b16):

I don't think that the inequalities would be connected on the graph. So one would be infinitely positive and the other infinitely negative

OpenStudy (crystal_b16):

If the X was in between both or 'and', then they'd be connected. Being that 'or is in there, then they'd go in the opposite direction of each other

OpenStudy (kaylieanime):

My hint with remembering with the direction inequalities is simple: the arrow points the direction it's headed. This applies to all inequality signs. \[> meaning \rightarrow on number line\] \[< meaning \leftarrow on number line\]